OpenStudy (anonymous):

Evaluate tan 30° without using a calculator by using ratios in a reference triangle.

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

Do you know what sin of 30 degrees is?

OpenStudy (anonymous):

And, do you know what cos of 30 degrees is?

OpenStudy (anonymous):

In radians or degrees?

OpenStudy (precal):

learn your special triangles first 45, 45, 90 30. 60. 90 60, 30, 90

OpenStudy (anonymous):

Ok I do know them

OpenStudy (anonymous):

Radians

OpenStudy (anonymous):

sin30= -0.9880316241

OpenStudy (anonymous):

cos30=0.1542514499

OpenStudy (anonymous):

No, exact values. There is no way you got thos values without a calculator. The question said not to use a calculator

OpenStudy (anonymous):

Ok no I don't know

OpenStudy (anonymous):

Oh wait yes I do

OpenStudy (anonymous):

sin30= 1/2

OpenStudy (anonymous):

Look at the unit circle

OpenStudy (anonymous):

cos30= \[\sqrt{3}/2\]

OpenStudy (anonymous):

To find tan, divide sin/ cos

OpenStudy (precal):

http://staffweb.psdschools.org/kemotich/Mrs_Motichka/Precalculus%20A/Chapter%205/Special%20Right%20Triangles%20and%20the%20Unit%20Circle.pdf look at this, this will help you learn the unit circle with the special rt triangles.

OpenStudy (anonymous):

Ok I got 1/\[\sqrt{3}\]

OpenStudy (precal):

http://www.youtube.com/watch?v=BnIfvgbC1fo if you want a video, here is one for you

OpenStudy (anonymous):

Is 1/sqrt3 correct?

OpenStudy (anonymous):

Is 1/sqrt3 correct?

OpenStudy (anonymous):

Thanks for everyone's help!